# Unveiling Challenges for LLMs in Enterprise Data Engineering

### v1

@@ 4. The Data Challenge
In this section, we quantify the anatomy of enterprise data by comparing real-world data from SAP databases to publicly available table corpora. We point out four challenges (C1-C4) specific to enterprise data and perform experiments to evaluate how they affect LLM performance. As an example task, we choose the task of column type annotation shown in Figure 2 (left).
@@ 4.1. Data Challenges
For this study, we constructed a new corpus SAPCTA . Table 2 compares the data characteristics of the real-world customer data in SAPCTA to several publicly available column type annotation benchmarks — WikiTables-TURL (Deng et al., 2020), SOTAB (Korini et al., 2022), GitTablesCTA (Hulsebos et al., 2021), and SportsTables (Langenecker et al., 2023). We observe the following differences:
C1: Table size. A first important observation is that enterprise tables typically have substantially more rows and columns than the tables in public corpora. As shown in Table 2, some tables have hundreds of columns and millions of rows. While the large scale is a well-established data management problem (Zhu et al., 2019), it poses challenges for LLMs, which have limited context windows. Although recent models have extended context windows, feeding large tables into LLMs still has downsides since latency and cost depend on the input size, and recent studies have shown that long contexts can lead to degraded performance for data residing in the “middle” (Liu et al., 2024a).
@@
C3: Sparsity. A third insight is that enterprise data is highly sparse. Table 2 shows that on average, 43% of the cells in enterprise tables are empty, compared to only 7-12% in existing datasets. This high level of sparsity results in a significant lack of information, which poses a challenge for LLMs that rely on contextual cues to make accurate predictions. Moreover, we find that in addition to empty values, the cells in enterprise tables often contain dummy values such as 00000 that also denote the absence of an actual value.
C4: Data types. Surprisingly, we find that only 45% of the columns in SAPCTA are classified as numerical data by pandas, challenging the common assumption that enterprise data is predominantly numerical (Langenecker et al., 2023). Moreover, we see that only 14% of columns in the database schema have numerical data types like INT and DECIMAL. A closer inspection of the actual data reveals that the non-numerical data type NVARCHAR is often used to store symbolic values and codes such as invoice and material numbers, which is in line with previous findings (Vogelsgesang et al., 2018). Since these values are not self-expressive, LLMs cannot make use of them without additional context.
@@ 4.2. Experiments & Results
To study how the challenges C1-C4 affect LLMs for data engineering, we compare the model performance on our real-world enterprise dataset SAPCTA to the performance on existing evaluation datasets.
Column type annotation task.
For the evaluation, we have chosen the task of column type annotation, as it is a well-established task where the goal is to annotate the columns of a relational table with semantic types from a pre-defined ontology (Hulsebos et al., 2019; Zhang et al., 2020a). We see it as an interesting example task to uncover the challenges of understanding enterprise tables with LLMs because it requires a semantic understanding of the content of each column as well as the values of other columns and the table and column names, which can all provide important signals to derive a semantic type.
@@
Our SAPCTA corpus spans diverse business domains such as Finance, Sales and Distribution, Material Management, and Production Planning. For our experiments, we select 100 representative tables from the larger corpus, which contains multiple thousands of tables. Our prompting strategy builds on best practices from existing literature, where the model annotates the columns of a given table based on a list of semantic types in the prompt and one randomly selected example (Zhang et al., 2020a; Korini and Bizer, 2023). Due to the LLMs’ limited context windows, we have to limit each table to three randomly selected rows. For similar reasons, we serialize the tables in CSV format, which requires fewer formatting tokens than other serialization schemes like Markdown and JSON (Singha et al., 2023; Sui et al., 2024). We instruct the model to generate the column types as a JSON-formatted list.
Exp. 1: Enterprise vs. public tables. In our first experiment, we compare the performance for column type annotation on our SAPCTA corpus with the performance on GitTablesCTA (Hulsebos et al., 2021) and SportsTables (Langenecker et al., 2023). We perform each experiment twice, with and without including the table and column names (i.e., the table schema) in the prompt. Existing evaluations typically leave out the column names, since the semantic types are directly derived from them and the task would thus become trivial. However, for SAPCTA , the task is much harder. Therefore, we want to investigate how much the additional information helps.
Table 3 shows the results of this experiment. We make the following key observations: (1) LLMs have severe issues with column type annotation on enterprise data, leading to substantially worse results compared to the web resources GitTablesCTA and SportsTables. Especially in the experiments without table and column names, the results on enterprise data are particularly poor (F1 scores of only up to 0.05), indicating that the enterprise data on its own contains few helpful signals. (2) Adding table and column names to the prompt improves the results on the enterprise data, but only up to 0.34 using Claude 3.5 Sonnet, which is still much lower than for web tables. The remaining performance gap could potentially be attributed to the non-descriptive schema, the extremely wide and sparse tables, and the complex data types described in Section 4.1.
@@
Exp. 2: Textual vs. numerical data. LLMs are known to often perform better on textual data than on numerical data (Frieder et al., 2023; Langenecker et al., 2023). To analyze this effect on enterprise data, we examine the performance for non-numerical and numerical columns in our SAPCTA corpus. Table 4 shows that, as expected, we see a higher performance on non-numerical enterprise data. By contrast, the results on the public benchmarks GitTablesCTA and SportsTables are inconclusive, with only small performance gaps between non-numerical and numerical columns that differ between models. The low performance on the SAPCTA dataset (F1 scores of only up to 0.41 for non-numerical and 0.27 for numerical data) indicates that numerical data in enterprise systems is even harder to understand than in web tables. Furthermore, the low scores for non-numerical columns stems from the fact that the enterprise tables often store identifiers like INV0014056 as type NVARCHAR, as explained in Section 4.1 (C4).
@@
Exp. 3: Table width and sparsity. To further investigate the performance gap between web tables and enterprise data, we incrementally adapt the enterprise tables to resemble the characteristics of web tables more closely. Since two of the main differences are table width and sparsity (see C1 and C3), we vary the number of columns per table by randomly sampling subsets of columns and vary the sparsity by initially selecting only non-sparse columns and randomly removing individual cell values. Figure 3 shows that increasing numbers of columns lead to substantially worse results, indicating that the large table widths in enterprise data are indeed a major problem for LLMs. Regarding sparsity, Figure 4 shows worse results with increased sparsity if no table and column names are provided, whereas with table and column names, increased sparsity does not change the results much. This indicates that for the enterprise data, LLMs rely primarily on the column headers to predict the semantic types and do not take the cell values into account.
@@ 4.3. Discussion
Our analysis of the representative customer data from SAP shows that enterprise data has fundamentally different characteristics, and our experiments demonstrate that these differences lead to substantial performance declines. We have seen that the performance decreases with higher sparsities and growing table sizes, indicating that LLMs are unreliable at scale and sometimes fail to uphold even the basic table structure by generating an incorrect number of labels. While we have shown this for the task of column type annotation, these data challenges will appear again throughout the following sections for other data engineering tasks like schema matching and entity matching, highlighting that they are, in fact, general and affect a broad range of tasks.
@@
We believe that overcoming these challenges requires better representations for enterprise data. A first promising direction is the development of foundation models for relational data (Vogel et al., 2022; Wu et al., 2025). While there is already a large body of research on foundation models for tabular data (Qu et al., 2025; He et al., 2025; Li et al., 2024; Yang et al., 2022; Eisenschlos et al., 2021; Herzig et al., 2020; Yin et al., 2020; Deng et al., 2020), computing representations for complex enterprise data is still an open research problem. Challenges include topics such as scalability and cost as well as complex data types and non-descriptive table and column names.
To mitigate technical problems such as context size limitations, we could for example apply windowing approaches to handle the large tables. However, this approach is clearly limited as important information may be lost, which is especially problematic due to the low self-expressiveness of the enterprise data. One way to alleviate this issue could be to better contextualize the raw data by integrating metadata such as data dictionaries, which contain textual descriptions for table and column names and for symbolic values. However, preparing such additional data for LLMs currently imposes high manual overheads, as we will discuss in Section 5.
@@ 5. The Task Challenge
A second dimension where data engineering is different in enterprise scenarios is the complexities of the tasks, as shown in Figure 2 (center). In particular, the academic definition of a task often makes several simplifying assumptions that do not hold in real-world scenarios. For example, when looking at entity matching, the assumption is typically made that we compare the similarity of individual rows of two tables (Peeters and Bizer, 2023; Konda et al., 2016). However, when looking at entity matching in enterprises, business entities often span across multiple tables, forming a graph of tuples. Moreover, enterprises typically approach data engineering with broad business goals in mind. Therefore, enterprise tasks are often composed of multiple individual steps and are thus often compounds of simpler tasks (e.g., combining schema matching and entity matching). In the following, we present a detailed discussion of how enterprise tasks differ in their nature from the complexity of tasks evaluated in academia.
@@ 5.1. Task Challenges
We base our analysis of task-specific challenges in enterprise data engineering (C5-C8) on the following two case studies.
@@
Entity matching is a task that often occurs in enterprise scenarios. In the following, we study a representative scenario of matching the bank statements of incoming payments to open invoices, which is shown in Figure 5. The scenario showcases many challenges in enterprise entity matching, as discussed below, including problems such as complex table structures and the fact that matching requires finding 1:N or even N:M matches.
@@
For our second case study, we look into another enterprise scenario where two databases from different companies must be integrated. To ensure a unified and consistent customer database, this process involves the individual steps of schema matching, customer record matching, and data integration, as illustrated in Figure 2 (center).
@@
In the following, we highlight the general challenges that arise in enterprise tasks and provide examples based on these two scenarios.
@@
C5: Entities span multiple tables. As mentioned before, entities in enterprise systems are often business objects that are represented by multiple rows stored in different connected tables. For example, in SAP systems, data pertaining to a particular material is scattered across the MARA (material type and basic statistics), MARC (manufacturing-related details), MBEW (valuation data), and other tables. The same holds true for our first case study (entity matching), which we later show in our evaluation. As such, for many data engineering tasks, including entity matching, one must either manually construct views that extract the fields relevant to the task into a single table, or approaches have to deal directly with the complex table structures that form a business entity.
@@
C6: Compound tasks. While data engineering tasks in research are usually addressed as isolated problems, such as deduplication (Papenbrock et al., 2015) and missing value imputation (Mei et al., 2021), tasks in enterprise contexts are typically approached on a more holistic level concerning broader business objectives. Instead of focusing on individual tasks, enterprises aim to solve end-to-end workflows, as highlighted by our second case study (data integration). While steps of end-to-end workflows can be executed sequentially, errors often propagate and amplify in later steps. As such, analyzing steps in isolation does not reveal the overall quality of the task in enterprise scenarios.
@@
C7: Task complexity. Moreover, even individual workflow steps can be more complex than their counterparts in the academic setting. In the following, we discuss these differences by the example of entity matching: (1) Matching between different types of entities: In public entity matching datasets, the entities to be matched are usually of the same type, such as e-commerce products, restaurants, and scholarly articles (Konda et al., 2016; Peeters et al., 2024). Meanwhile, enterprise scenarios often require matching between different types of entities, like products to commodity codes or payments to invoices (as in our first case study), which have overlapping but different sets of attributes, as shown in Figure 5. (2) Multi-matches: An additional challenge in enterprise scenarios is that the matches are often not 1:1 matches as in the literature but can also be 1:N, N:1, or even N:M matches, making the problem much more difficult. In our payment-to-invoice matching scenario, it is quite common for a customer to pay multiple invoices with only one payment (1:N), or for one invoice to be paid by multiple payments (N:1), such as in the case of down payments or by holding back money due to quality issues with the product. Even a combination of both cases is possible (N:M).
@@
C8: Data complexity on top. Along with these task complexities, the complexity of the data makes the tasks themselves also harder. For example, when matching payments to invoices (first case study), the memo lines of incoming bank statements are not standardized, but rather free-form text entered by customers. Therefore, complex errors occur regularly, making automated matching difficult. Compared to such data errors in enterprise scenarios, the data errors in academic entity matching datasets are often much simpler.

### v2

@@ 4. The Data Challenge
In this section, we quantify the anatomy of enterprise data by comparing real-world customer data from SAP to publicly available table corpora. We point out four challenges specific to enterprise data and perform experiments to evaluate how they affect LLMs, using column type annotation as an example task (Figure 2 left).
@@ 4.1. Data Challenges
For our study, we constructed a new corpus called SAPCTA . Table 2 compares the data characteristics of the real-world customer data in SAPCTA to several publicly available column type annotation benchmarks: WikiTables-TURL (Deng et al., 2020), SOTAB (Korini et al., 2022), GitTablesCTA (Hulsebos et al., 2021), and SportsTables (Langenecker et al., 2023). We observe the following differences:
C1: Table size. A first important observation is that enterprise tables often have substantially more rows and columns than the tables in public corpora. As shown in Table 2, some tables have hundreds of columns and millions of rows. While the large scale is a well-established data management problem (Zhu et al., 2019), it poses challenges for LLMs, which have limited context windows. Although recent models have extended context windows, feeding large tables into LLMs still has downsides since latency and cost depend on the input size, and recent studies have shown that long contexts can lead to degraded performance for data residing in the “middle” (Liu et al., 2024).
@@
C3: Sparsity. A third insight is that enterprise data is highly sparse. Table 2 shows that on average, 43% of the cells in enterprise tables are empty, compared to only 7-12% in existing datasets. The high sparsity results in a significant lack of information, posing a challenge for LLMs which rely on contextual cues to make accurate predictions. Moreover, we find that in addition to empty values, the cells in enterprise tables often contain dummy values such as 00000 that also denote the absence of an actual value.
C4: Data types. Interestingly, we find that only 45% of the columns in SAPCTA are classified as numerical by pandas, suggesting that enterprise datasets can be more text-heavy than found in previous studies (Langenecker et al., 2023). Moreover, only 14% of the columns in the database schema have numerical data types like INT and DECIMAL. A closer inspection of the actual data reveals that the non-numerical data type NVARCHAR is often used to store symbolic values and codes such as invoice and material numbers, which is in line with previous findings (Vogelsgesang et al., 2018). Since these values are not self-expressive, LLMs cannot make use of them without additional context.
@@ 4.2. Experiments & Results
To study how the challenges C1-C4 affect LLMs for data engineering, we compare the models’ performance on our real-world enterprise dataset SAPCTA to their performance on existing evaluation datasets.
Task.
We evaluate on column type annotation, a well-established data engineering task where the goal is to annotate the columns of a relational table with semantic types from a pre-defined ontology (Hulsebos et al., 2019; Zhang et al., 2020). We see it as an interesting example task to uncover the challenges of understanding enterprise tables with LLMs, since it requires a semantic understanding of the content of each column as well as the values of other columns and the table schema, all of which provide important signals to derive a semantic type.
@@
Our SAPCTA corpus spans diverse business domains such as Finance, Sales and Distribution, Material Management, and Production Planning. For our experiments, we select 100 representative tables from the larger corpus, which contains multiple thousands of tables. The selection is based on discussions with SAP experts to identify the most widely used tables in the system. In total, the dataset includes 5,089 semantic types, such as Amount Difference in Local Currency or Product Cost Collector, which are generally more fine-granular than those used in other datasets (Korini et al., 2022; Hulsebos et al., 2021).
Our prompting strategy builds on best practices from existing literature, where the model annotates the columns of a given table based on a list of semantic types and one randomly-selected example included in the prompt (Zhang et al., 2020; Korini and Bizer, 2023).222To focus solely on the data challenges, we include only a subset of the semantic types in every prompt similar to the setup from Figure 1 (SAPCTA ). Based on ablations (see Appendix A.1 in our extended technical report), we limit each table to three randomly-selected rows as a reasonable trade-off between performance and cost. For similar reasons, we serialize the tables in CSV format, which requires fewer formatting tokens than other serialization schemes like Markdown and JSON (Singha et al., 2023; Sui et al., 2024). We instruct the model to generate the column types as a JSON-formatted list.
Exp. 1: Enterprise vs. public tables. In our first experiment, we compare the performance for column type annotation on our SAPCTA corpus with the performance on GitTablesCTA (Hulsebos et al., 2021) and SportsTables (Langenecker et al., 2023). We perform each experiment twice, with and without including the table and column names (i.e., the table schema) in the prompt. Existing evaluations typically leave out the column names, since the semantic types are directly derived from them and the task would thus become trivial. For SAPCTA , however, the task is much harder. Therefore, we want to investigate how much the additional information helps.
Table 3 shows the results of this experiment. We make the following observations: First, LLMs have severe issues with column type annotation on enterprise data, leading to substantially worse results compared to the public benchmarks GitTablesCTA and SportsTables. The results are particularly poor in the experiments without table and column names, indicating that enterprise data on its own contains few helpful signals. Adding table and column names to the prompt improves the results, but they still remain much lower than for web tables. The remaining performance gap could potentially be attributed to the non-descriptive schema, the extremely wide and sparse tables, and the complex data types.
@@
Exp. 2: Textual vs. numerical data. LLMs are known to often perform better on textual data than on numerical data (Frieder et al., 2023; Langenecker et al., 2023). To study this effect on enterprise data, we compare the performance for non-numerical and numerical columns in our SAPCTA corpus.
Table 4 shows that, as expected, we see a higher performance on non-numerical enterprise data. By contrast, the results on the public benchmarks GitTablesCTA and SportsTables remain inconclusive, with only small performance gaps between non-numerical and numerical columns that also differ between models. The low performance on the SAPCTA dataset indicates that numerical data in enterprise systems is even harder to understand than in public benchmarks. Furthermore, the low scores for non-numerical columns may stem from the fact that enterprise tables often store identifiers like INV0014056 as type NVARCHAR.
Exp. 3: Table width and sparsity. To further investigate the performance gap between public benchmarks and enterprise data, we incrementally adapt the enterprise tables to resemble the characteristics of web tables more closely. Since two of the main differences are table width and sparsity, we vary the number of columns per table by randomly sampling subsets of columns and vary the sparsity by initially selecting only non-sparse columns and randomly removing individual cell values.
Figure 3 shows that increasing table widths lead to substantially worse results, indicating that the large table widths in enterprise data are indeed a major problem for LLMs. Figure 4 shows that increased sparsity leads to worse results only if no table and column names are provided, whereas with table and column names, increased sparsity does not change the results much. This indicates that on enterprise data, LLMs rely primarily on the column names to predict the semantic types, maintaining almost constant accuracy even if no cell values are provided (Figure 4 left, sparsity of 1.0).
@@
Exp. 4: Improvement strategies. Figure 3 shows that column type annotation accuracy decrases as the table width increases. To mitigate this effect, we explore three strategies for handling wide tables: All columns (full table) is our baseline setup, where the model predicts all semantic types of the full table in a single call. Chunks of columns divides the table into chunks of columns and prompts the model for each chunk separately without the full table as context. One column at a time (with full table context) predicts each column type in a separate LLM call, but provides the full table in the prompt. We state the column to be predicted either by name or by index to analyze which variant works better.
Table 5 shows that on public benchmarks, the differences between strategies are small, although smaller chunk sizes reduce performance if no column names are provided. This indicates that the data in public benchmarks provides strong contextual signals for column type annotation. For SAPCTA , however, chunking without column names does not further degrade the already low performance. With column names, chunking even improves results. Notably, for SAPCTA , the one-column-at-a-time strategy using the full table as context yields the best results. Since enterprise data without headers provides little signal, we suspect that the models rely primarily on the column names, largely ignoring the actual data. This method is also the most expensive, requiring one LLM call per column with the full table as context. Inference over the full SAPCTA dataset with GPT-4o costs USD 53.85, compared to USD 4.37 for the baseline—an order of magnitude more. Yet even with this setup, performance remains far below that on public benchmarks.
@@ 4.3. Discussion
Our analysis of the customer data from SAP shows that enterprise data differs fundamentally from public benchmarks, and our experiments demonstrate that these differences lead to substantial performance decreases. The performance declines with higher sparsities and larger table widths, indicating that LLMs are unreliable at scale and sometimes fail to uphold even the basic table structure by generating an incorrect number of labels. Moreover, the non-descriptive schemas and complex data types pose significant challenges for LLMs. While we have shown this for the task of column type annotation, these data challenges will appear again throughout the following sections for other data engineering tasks, highlighting that they are general and affect a broad range of tasks.
@@
We believe that overcoming these challenges requires better representations for enterprise data. A first promising direction is the development of foundation models for relational data (Vogel et al., 2022; Wu et al., 2025). While there is already a large body of research on foundation models for tabular data (Qu et al., 2025; He et al., 2025; Li et al., 2024; Yang et al., 2022; Eisenschlos et al., 2021; Herzig et al., 2020; Yin et al., 2020; Deng et al., 2020), computing representations for complex enterprise data is still an open research problem. Some LLM weaknesses like their low reliability at scale require rethinking how they are applied to enterprise data, for example by using chunking to handle large tables. One way to alleviate the low self-expressiveness of the enterprise data could be to better contextualize the raw data by incorporating metadata such as data dictionaries, which contain textual descriptions for table and column names and for symbolic values.
| GitTablesCTA | SportsTables | SAPCTA | |||||
| Column Names | w/out | with | w/out | with | w/out | with | |
| GPT-4o-Mini | All columns (full table) | 0.52 | |||||
| Chunks of 50 columns | 0.52 | 0.29 | |||||
| Chunks of 10 columns | 0.97 | 0.62 | |||||
| Chunks of 1 columns | |||||||
| 1 column by index (full table) | 0.03 | ||||||
| 1 column by name (full table) | - | - | - | 0.32 | |||
| GPT-4o | All columns (full table) | 0.56 | 0.99 | 0.91 | |||
| Chunks of 50 columns | 0.99 | 0.58 | |||||
| Chunks of 10 columns | 0.99 | 0.91 | |||||
| Chunks of 1 columns | |||||||
| 1 column by index (full table) | 0.05 | ||||||
| 1 column by name (full table) | - | - | - | 0.52 | |||
@@ 5. The Task Challenge
A second dimension where data engineering in enterprise settings stands out is the complexity of the tasks. The academic definition of a data engineering task often makes simplifying assumptions that do not hold in real-world scenarios. For example, literature on entity matching often assumes that we compare the similarity of individual rows of two tables (Peeters and Bizer, 2023; Konda et al., 2016). By contrast, business entities in enterprises often span across multiple tables, forming graphs of tuples. At the same time, enterprises approach data engineering with broad business objectives in mind. Therefore, enterprise tasks are often composed of multiple steps (e.g., combining schema matching and entity matching). In this section, we discuss how enterprise tasks differ in their nature from the tasks evaluated in academia and examine how these differences affect LLMs.
@@ 5.1. Task Challenges
We base our analysis of the task-related challenges in enterprise data engineering on the following two case studies:
@@
Entity matching is a data engineering task that often occurs in enterprise scenarios. We study the scenario shown in Figure 5, in which the bank statements of incoming payments must be matched to a company’s open invoices. The scenario showcases many challenges of entity matching in enterprises, including problems like complex table structures and the fact that matching requires finding 1:N or even N:M matches.
@@
For our second case study, we look into another enterprise scenario where two customer databases from different companies must be merged. To ensure a unified and consistent customer database, this process involves the individual steps of schema matching, customer record matching, and data integration, as illustrated in Figure 2 (center).
@@
In the following, we highlight the general challenges that arise in enterprise tasks and provide examples based on the two scenarios.
@@
C5: Entities span multiple tables. Entities in enterprise systems are often business objects represented by multiple rows stored in different connected tables. In SAP systems, data pertaining to a particular material is scattered across the MARA (material type and basic statistics), MARC (manufacturing-related details), MBEW (valuation data), and other tables. The same holds true for the entities in our first case study (entity matching). For many data engineering tasks, one must therefore either manually construct views that extract the relevant fields into a single table, or approaches must work directly with the complex table structures that form a business entity.
@@
C6: Compound tasks. Whereas data engineering in research is often addressed as isolated problems, such as deduplication (Papenbrock et al., 2015) and missing value imputation (Mei et al., 2021), tasks in enterprise contexts are typically approached on a more holistic level concerning broader business objectives. Instead of focusing on individual tasks, enterprises aim to solve end-to-end workflows, as highlighted by our second case study (database integration). While the steps of such compound tasks could be executed sequentially, errors often propagate and amplify in later steps. Analyzing each step in isolation does not reveal the overall quality of the task in enterprise scenarios.
@@
C7: Task complexity. Even individual workflow steps can be more complex than their counterparts studied in academia. For example, the entities to be matched in public entity matching datasets are usually of the same type, such as e-commerce products, restaurants, and scholarly articles (Konda et al., 2016; Peeters et al., 2024). By contrast, enterprise scenarios often require matching between different types of entities, like products to commodity codes or payments to invoices, which have overlapping but different sets of attributes. An additional challenge in enterprise scenarios is that the matches are often not 1:1 matches as in the literature, but can also be 1:N, N:1, or even N:M matches, making the problem much harder. In our payment-to-invoice matching scenario, it is common for a customer to pay multiple invoices with only one payment (1:N), or for one invoice to be paid by multiple payments (N:1), such as in the case of down payments.
@@
C8: Data complexity on top. Along with these task complexities, the complexity of the data makes the tasks themselves also harder. For example, when matching payments to invoices, the memo lines of incoming bank statements are not standardized, but rather free-form text entered by customers. Therefore, complex errors occur regularly, making automated matching difficult. Compared to such data errors in enterprise scenarios, the errors in academic entity matching datasets are often much simpler.